OpenStudy (anonymous):

What is the solution to the equation? -8x + 9x - 21 = -12 - 16 { }

OpenStudy (anonymous):

-12 - 16 = -28 -8 + 9 - 21 = -20 X = ?

OpenStudy (anonymous):

not -2 that would = 23

OpenStudy (anonymous):

-8 is also a no

OpenStudy (anonymous):

@Nnesha

Nnesha (nnesha):

\(\color{blue}{\text{Originally Posted by}}\) @Demonx341 -12 - 16 = -28 -8 + 9 - 21 = -20 X = ? \(\color{blue}{\text{End of Quote}}\) you can't `can't` combine -9and -8 with -21

OpenStudy (anonymous):

? its -8 + 9 - 21

Nnesha (nnesha):

they are not like terms you can't combine variable term with just number

OpenStudy (anonymous):

oh XD

Nnesha (nnesha):

it's -8x+9x-21 combine like terms here is an example 2x+2x +5= 4x+5 i can't combine 4x with 5

OpenStudy (ryanwhite499):

Let's solve your equation step-by-step. −8x+9x−21=−12−16 Step 1: Simplify both sides of the equation. −8x+9x−21=−12−16 Simplify: (Show steps) x−21=−28 Step 2: Add 21 to both sides. x−21+21=−28+21 x=−7 Answer: x=−7

OpenStudy (anonymous):

-8x + 9x = 1x

Nnesha (nnesha):

let him do it.

OpenStudy (anonymous):

btw i have to find out what x = XD

OpenStudy (anonymous):

It would Be -7 Tho :P

Nnesha (nnesha):

yes right x -21=-28 right now move the 21 to the right side

OpenStudy (ryanwhite499):

Yes... In my equation... x = -7

OpenStudy (anonymous):

ok XD

OpenStudy (ryanwhite499):

Great Work!

OpenStudy (anonymous):

-28 - 21 = -49

Nnesha (nnesha):

nope how would you cancel out -21 from left side would you add ,subtract ,divide or multiply ?

OpenStudy (anonymous):

add

OpenStudy (anonymous):

-28 + 21= -7

Nnesha (nnesha):

looks good!

OpenStudy (anonymous):

1*-7 = -7 xD

Nnesha (nnesha):

now you can check ur answer substitute x for -7 \[\huge\rm -8(-7)+9(-7)-21=-12-16\] both sides are equal good to go!

OpenStudy (anonymous):

:P both = -28

Nnesha (nnesha):

yes right that means x=-7 is correct good work!

OpenStudy (ryanwhite499):

Demon you need to fan me in order for me to send messages back

OpenStudy (anonymous):

XD forgot about that

OpenStudy (ryanwhite499):

Sorry. Please continue